OpenStudy (anonymous):

Write the equation of the line that has a slope of -3/4 and passes through (-2,1) a. y = -3/4x + 5/2 b. y = -3/4 x - 1/2 c. y = -3/4x - 5/4 d. y = -3/4x - 11/4

OpenStudy (aroub):

y=mx+b 1=-3/4(-2)+b 1=3/2+b -3/2 +2/2=b -1/2=b y=-3/4x-1/2 so the answer is b.

OpenStudy (anonymous):

Thank yu! Yur so helpful :))

OpenStudy (aroub):

no prob! anytime! :D